the asteroid Ceres was recently reclassified as a dwarf planet by the International Astronomical Union. It has a mass of 9.43 * 10^20 kg, and a radius of about 470, THe acceleration due to gravity on the surface of Ceres is ....... m/s^2
1
Expert's answer
2020-06-09T13:20:19-0400

We can find the acceleration due to gravity on the surface of Ceres from the formula:


"g = \dfrac{GM}{r^2},""g = \dfrac{6.67 \cdot 10^{-11} \ \dfrac{N \cdot m^2}{kg^2} \cdot 9.43 \cdot 10^{20} \ kg}{(4.7 \cdot 10^5 \ m)^2} = 0.28 \ \dfrac{m}{s^2}."

Answer:

"g = 0.28 \ \dfrac{m}{s^2}."